How many bit strings of length 14 contains the following types

(I) atmost three is?

(ii) an equal number of 0s and 1s

(iii) an odd number of 1s



1
Expert's answer
2021-03-30T10:54:01-0400

I. Atmost three

"C(n,r)=\frac{n!}{(n-r)!r!}\\
n=14\\
r≤3\\
=C(14,0)+C(14,1)+C(14,2)+C(14,3)\\
=1+14+91+364=470"

II. An equal number of 0s and 1s. This means that there will be 7 1s and 7 0s.

"n=14\\
r=7\\
=C(14,7)\\
=3432"

III. An odd number of 1s. This means that 1s can be 1,3,5,7,9,11,13.


"=C(14,1)+C(14,3)+C(14,5)+C(14,7)+C(14,9)+C(14,11)+C(14,13)\\
=14+364+2002+3432+2002+364+14\\
=8192"